OpenStudy (anonymous):

The Miller and Benson families went to a theme park. The Millers bought 6 adult and 15 child tickets for $423. The Bensons bought 5 adult and 9 child tickets for $293. Find the cost of each type of ticket.

OpenStudy (calculusxy):

x=adults y=children Finding the y value: -5(6x+15y=$423)---> -30x-25y=-215 6(5x+9y=$293)---->. 30x-54y=1758 (cross out -30x and 30 because. ------------------ -21y=357 and y=17 they are like terms) Finding the x: 6x+15(17)=423 6x+225=423 6x=168 x=28 So y=17 x=28

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."